# AGZO: Activation-Guided Zeroth-Order Optimization for LLM Fine-Tuning

### v1

@@ A.3 AGZO defeat MEZO in cosine similarity
$\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0}^{\text{AGZO}},G\big)\right]=\beta_{mr}\,\frac{\|GP_{A}\|_{F}}{\|G\|_{F}}\!,\qquad\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0}^{\text{MEZO}},G\big)\right]=\beta_{mn},$
where $A\in\mathbb{R}^{n\times r}$ is orthonormal (AGZO’s subspace), $P_{A}=AA^{\top}$, and
@@
$\mathbb{E}_{R}\!\left[\cos(\widehat{G}_{0}^{\text{AGZO}},G)\right]\;>\;\mathbb{E}_{R}\!\left[\cos(\widehat{G}_{0}^{\text{MEZO}},G)\right]\quad\Longleftrightarrow\quad\alpha(A;G)\;>\;\frac{\beta_{mn}}{\beta_{mr}}.$ (52)
@@
$\frac{\beta_{mn}}{\beta_{mr}}=\frac{\Gamma(\frac{mn}{2})}{\Gamma(\frac{mr}{2})}\cdot\frac{\Gamma(\frac{mr+1}{2})}{\Gamma(\frac{mn+1}{2})}.$
By lemma 5.6, we have
$\frac{\beta_{mn}}{\beta_{mr}}<\frac{\sqrt{\frac{mr}{2}}}{\sqrt{\frac{mn}{2}-\frac{1}{2}}}=\frac{\sqrt{mr}}{\sqrt{mn-1}}.$
@@
$\alpha(A;G)\;>\;\frac{\sqrt{mr}}{\sqrt{mn-1}}$
or (by equation (31))
$\frac{\sum_{i=1}^{r}B_{ii}\,\sigma_{i}^{2}}{\sum_{i=1}^{s}B_{ii}\,\sigma_{i}^{2}}>\frac{r}{n-1/m}$ (53)
@@
Consider a layer with gradient factorization $\nabla_{W_{\ell}}F(W)=Q_{\ell}H_{\ell}^{\top}$ and compact SVD $H_{\ell}=U_{\ell}\Sigma_{\ell}V_{\ell}^{\top}$ of rank $s_{\ell}<n_{\ell}$ with singular values $\{\sigma_{\ell,i}\}_{i=1}^{s_{\ell}}$. Let $A_{\ell}=U_{\ell}^{(r_{\ell})}$ be the AGZO subspace and $B_{\ell}=V_{\ell}^{\top}Q_{\ell}^{\top}Q_{\ell}V_{\ell}$. If the average of its first $r$ diagonal entries is not less than the average of all diagonal entries, i.e.,
$\frac{1}{r_{\ell}}\sum_{i=1}^{r_{\ell}}B_{\ell,ii}\;\geq\;\frac{1}{s_{\ell}}\sum_{i=1}^{s_{\ell}}B_{\ell,ii},$ (54)
and that $H_{\ell}$ is low-rank, i.e., $s_{\ell}<n_{\ell}$. Then
$\frac{\sum_{i=1}^{r_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}{\sum_{i=1}^{s_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}\;>\;\frac{r_{\ell}}{s_{\ell}}\;\geq\;\frac{r_{\ell}}{n_{\ell}-1/m_{\ell}},$ (55)
@@
$\frac{\sum_{i=1}^{r}B_{ii}\,\sigma_{i}^{2}}{\sum_{i=1}^{s}B_{ii}\,\sigma_{i}^{2}}>\frac{r}{s}\geq\frac{r}{n-1/m}.$
@@
The second inequality is due to the fact that $s<n$ and both $s,n\in\mathbb{N}^{+}$ . Combining with equation (53) and (52), we can conclude that AGZO generally can get gradient estimation have larger cosine similarity with the true gradient than MEZO.
Appendix B Detail of Pangu Model Experiments
This section presents our additional experimental results of Pangu-1B model on Ascend NPU.
Model.
To evaluate the generality of our approach across different model architectures, we further conduct experiments on the openPangu family of models. The openPangu series has demonstrated strong performance on a wide range of general language understanding and sophisticated reasoning tasks (Rang et al., 2025; Chen et al., 2025a). In particular, we select openPangu-embedded-1B [^1] https://huggingface.co/FreedomIntelligence/openPangu-Embedded-1B, a model specifically designed for efficient inference on edge devices (we use the GPU variant in our experiments [^2] https://modelscope.cn/models/wangrongsheng/openPangu-Embedded-1B-model-GPU).
Evaluation Setup.
Under the same experimental setup as described in the main pages, we conduct evaluations on openPangu-embedded-1B using SuperGLUE (Wang et al., 2019) and additional datasets, with identical baseline comparisons. All training experiments are conducted on a server running Ubuntu 22.04, equipped with 96GB RAM and two NVIDIA RTX 3090 GPUs. For inference evaluation, in addition to the GPU platform described above, we also perform evaluations on a high-memory NPU server configured with 1.5 TB of RAM and 8 $\times$ 64 GB 910B2 NPUs running EulerOS 2.0.
Training Details.
@@ Appendix B Detail of Pangu Model Experiments
Regarding numerical precision, unlike the main experiments where the Qwen models are loaded in FP32 precision, we load the openPangu models using BF16 precision. Since FP32 has 16 more mantissa bits than BF16, zeroth-order optimization requires a perturbation magnitude that is 2–3 orders of magnitude larger when using BF16 compared to FP32. Accordingly, for the openPangu models, we set the zeroth-order perturbation parameter to $\textsc{zo\_eps}=1\times 10^{-4}$. We adopt the same experimental configurations, i.e., training, preprocessing, evaluation, and hyperparameter tuning setup as described in Section 6.
B.1 End-to-end cross-platform fine-tuning performance
We evaluate the cross-platform inference performance, i.e, evaluating the performance on NPU with GPU-trained models. Table 3 presents the performance of AGZO and various baselines on openPangu-embedded-1B across on NPU. Across both GPU and NPU, AGZO consistently achieves the best results among zeroth-order methods and non-training baselines on most downstream tasks. On the NPU, AGZO attains an average score of 0.709, outperforming other ZO baselines on tasks such as SST2, COPA, BoolQ, RTE, and WiC. The slightly lower performance on the NPU compared to the GPU may be attributed to subtle differences in numerical precision, memory layout, or low-level kernel implementations, which can affect the propagation of small perturbations used in zeroth-order optimization.
Task FO AGZO MEZO LOZO Zero ICL
SST2 0.821 0.765 0.766 0.718 0.571 0.710
COPA 0.800 0.770 0.740 0.720 0.760 0.740
CB 0.696 0.696 0.732 0.643 0.482 0.446
BoolQ 0.752 0.728 0.697 0.694 0.696 0.731
RTE 0.780 0.729 0.729 0.682 0.578 0.682
WiC 0.657 0.567 0.552 0.542 0.469 0.495
Avg. 0.738 0.709 0.703 0.667 0.593 0.636
B.2 Peak GPU memory footprint
@@
We empirically validate our memory efficiency analysis by measuring the peak GPU memory footprint of fine-tuning Pangu-1B on the DROP task. The comparative results are presented in Figure 4. In Figure 4(a), we fix the batch size at 4 and scale the sequence length. While the First-Order (FO) baseline (SGD) encounters Out-Of-Memory (OOM) errors at a sequence length of 1,536, AGZO maintains a strictly linear scaling, consuming only 9.69 GB even at a length of 2,048.
@@
Moreover, in the second setting with a fixed sequence length of 256 (Figure 4(b)), SGD triggers OOM at a batch size of 32, whereas AGZO remains operative up to a batch size of 64 ($\approx$15.15 GB).
The experimental results demonstrate that the memory trajectory of AGZO aligns almost perfectly with other forward-only zero-order baselines (i.e., MEZO and LOZO). This evidence substantiates that the additional computation required for our activation-guided subspace construction introduces negligible memory overhead, preserving the inherent memory advantages of ZO optimization.

### v2

@@ A.3 AGZO defeat MEZO in cosine similarity
$\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0}^{\text{AGZO}},G\big)\right]=\beta_{d_{out}r}\,\frac{\|GP_{A}\|_{F}}{\|G\|_{F}}\!,\qquad\mathbb{E}_{R}\!\left[\cos\!\big(\widehat{G}_{0}^{\text{MEZO}},G\big)\right]=\beta_{d_{out}d_{in}},$
where $A\in\mathbb{R}^{d_{in}\times r}$ is orthonormal (AGZO’s subspace), $P_{A}=AA^{\top}$, and
@@
$\mathbb{E}_{R}\!\left[\cos(\widehat{G}_{0}^{\text{AGZO}},G)\right]\;>\;\mathbb{E}_{R}\!\left[\cos(\widehat{G}_{0}^{\text{MEZO}},G)\right]\quad\Longleftrightarrow\quad\alpha(A;G)\;>\;\frac{\beta_{d_{out}d_{in}}}{\beta_{d_{out}r}}.$ (43)
@@
$\frac{\beta_{d_{out}d_{in}}}{\beta_{d_{out}r}}=\frac{\Gamma(\frac{d_{out}d_{in}}{2})}{\Gamma(\frac{d_{out}r}{2})}\cdot\frac{\Gamma(\frac{d_{out}r+1}{2})}{\Gamma(\frac{d_{out}d_{in}+1}{2})}.$
By lemma A.7, we have
$\frac{\beta_{d_{out}d_{in}}}{\beta_{d_{out}r}}<\frac{\sqrt{\frac{d_{out}r}{2}}}{\sqrt{\frac{d_{out}d_{in}}{2}-\frac{1}{2}}}=\frac{\sqrt{d_{out}r}}{\sqrt{d_{out}d_{in}-1}}.$
@@
$\alpha(A;G)\;>\;\frac{\sqrt{d_{out}r}}{\sqrt{d_{out}d_{in}-1}}$
or (by taking square)
$\frac{\sum_{i=1}^{r}B_{ii}\,\sigma_{i}^{2}}{\sum_{i=1}^{s}B_{ii}\,\sigma_{i}^{2}}>\frac{r}{d_{in}-1/d_{out}}$ (44)
@@
Consider a layer with gradient factorization $\nabla_{W_{\ell}}F(W)=Q_{\ell}H_{\ell}^{\top}$ and compact SVD $H_{\ell}=U_{\ell}\Sigma_{\ell}V_{\ell}^{\top}$ of rank $s_{\ell}<d_{in}$ with singular values $\{\sigma_{\ell,i}\}_{i=1}^{s_{\ell}}$. Let $A_{\ell}=U_{\ell}^{(r)}$ be the AGZO subspace and $B_{\ell}=V_{\ell}^{\top}Q_{\ell}^{\top}Q_{\ell}V_{\ell}$. If the average of its first $r$ diagonal entries is not less than the average of all diagonal entries, i.e.,
$\frac{1}{r}\sum_{i=1}^{r}B_{\ell,ii}\;\geq\;\frac{1}{s_{\ell}}\sum_{i=1}^{s_{\ell}}B_{\ell,ii},$ (45)
and that $H_{\ell}$ is low-rank, i.e., $s_{\ell}<d_{in}$. Then
$\frac{\sum_{i=1}^{r}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}{\sum_{i=1}^{s_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}\;>\;\frac{r}{s_{\ell}}\;\geq\;\frac{r}{d_{in}-1/d_{out}},$ (46)
@@
$\frac{\sum_{i=1}^{r}B_{ii}\,\sigma_{i}^{2}}{\sum_{i=1}^{s}B_{ii}\,\sigma_{i}^{2}}>\frac{r}{s}\geq\frac{r}{d_{in}-1/d_{out}}.$
@@
The second inequality is due to the fact that $s<d_{in}$ and both $s,d_{in}\in\mathbb{N}^{+}$ . Combining with equation (44) and (43), we can conclude that AGZO generally can get gradient estimation have larger cosine similarity with the true gradient than MEZO.
Appendix B Experimental Details
B.1 Datasets
We evaluate our method on a diverse set of natural language understanding and question-answering benchmarks. The datasets used in our experiments include:
  • SST-2 (Socher et al., 2013): A single-sentence classification task focusing on sentiment analysis of movie reviews.
  • DROP (Dua et al., 2019): A reading comprehension benchmark that requires discrete reasoning (e.g., sorting, counting, arithmetic) over paragraphs to answer questions. This dataset challenges the model’s ability to perform complex logical operations beyond simple span extraction.
  • SuperGLUE Benchmark Tasks (Wang et al., 2019): We select several challenging tasks requiring complex reasoning:
    • BoolQ (Clark et al., 2019): QA task with yes/no answers based on passages.
    • MultiRC (Khashabi et al., 2018): QA task requiring reasoning over multiple sentences.
    • RTE: Natural language inference task.
    • WiC (Pilehvar & Camacho-Collados, 2019): Word sense disambiguation task.
    • CB (De Marneffe et al., 2019): Few-shot multi-class entailment task.
    • COPA (Roemmele et al., 2011): Causal reasoning task.
  • SQuAD (Rajpurkar et al., 2016): A standard reading comprehension dataset based on Wikipedia articles.
B.2 Implementation and Hyperparameters
All zeroth-order optimization methods (AGZO, MeZO, and LOZO) are implemented using the same codebase to ensure a fair comparison. For all ZO experiments, we perform fine-tuning for a total of 20,000 steps. This fixed budget allows us to directly compare the convergence speed and final performance of different estimators under identical computational constraints. We set the smoothing parameter (perturbation scale) $\mu=10^{-7}$ for all methods (AGZO, MeZO, and LOZO) on Qwen3 models and $\mu=10^{-4}$ on the Pangu model. This value was chosen to minimize discretization error while maintaining numerical stability.
For AGZO, we set the subspace rank $r_{\ell}=1$ for all linear layers, a design choice driven by both signal quality and memory efficiency. As shown in our spectral analysis (Figure 1), the singular values of activation matrices decay rapidly; thus, a rank-1 setting concentrates the finite-difference perturbation along the single most dominant direction of the activation landscape. This maximizes the signal-to-noise ratio of the gradient estimate by avoiding the exploration of low-energy directions that contribute little to the true gradient. Furthermore, this rank-1 basis minimizes the memory overhead for storing the subspace information ($A_{\ell}$), ensuring the peak memory footprint remains nearly identical to that of standard inference.
B.3 Additional Experiments on Pangu-1B
In this section, we provide supplementary experimental details for the openPangu-embedded-1B model (Chen et al., 2025a). While the main text (Section 6) reports the standard GPU fine-tuning performance, here we focus on (1) specific implementation details regarding low-precision (BF16) optimization, (2) cross-platform verification on Ascend NPUs, and (3) memory footprint analysis specific to this architecture.
Precision and Hyperparameters (BF16).
@@ B.3 Additional Experiments on Pangu-1B
Unlike the Qwen experiments where models are loaded in FP32, we conduct Pangu experiments using BF16 precision to simulate resource-constrained edge scenarios. Since FP32 has significantly higher mantissa precision than BF16, zeroth-order optimization requires a larger perturbation magnitude to overcome numerical noise. Accordingly, we adjust the perturbation parameter to $\mu=1\times 10^{-4}$ for Pangu (compared to $10^{-7}$ for FP32 models).
Peak Memory Footprint on Pangu-1B.
@@
We empirically validate the memory efficiency of AGZO on the Pangu architecture by measuring the peak GPU memory footprint during fine-tuning on the DROP task with Testbed One. As shown in Figure 4(a), when fixing the batch size at 4 and scaling the sequence length, the First-Order (FO) baseline encounters Out-Of-Memory (OOM) errors at a sequence length of 1,536. In contrast, AGZO maintains a strictly linear scaling, consuming only 9.69 GB even at a context length of 2,048. Similarly, with a fixed sequence length of 256 (Figure 4(b)), SGD triggers OOM at a batch size of 32, whereas AGZO remains operative up to a batch size of 64. This confirms that the memory advantages of AGZO observed on Qwen models (Section 6.4) generalize to different architectures, with the activation-guided subspace construction introducing negligible memory overhead.